# Dataset Distillation of 3D Point Clouds via Distribution Matching

### v2

@@ Abstract
Large-scale datasets are usually required to train deep neural networks, but it increases the computational complexity hindering the practical applications. Recently, dataset distillation for images and texts has been attracting a lot of attention, that reduces the original dataset to a synthetic dataset to alleviate the computational burden of training while preserving essential task-relevant information. However, the dataset distillation for 3D point clouds remains largely unexplored, as the point clouds exhibit fundamentally different characteristics from that of images, making the dataset distillation more challenging. In this paper, we propose a distribution matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures as well as the orientations of the synthetic 3D objects. To address the semantic misalignment caused by unordered indexing of points, we introduce a Semantically Aligned Distribution Matching loss computed on the sorted features in each channel. Moreover, to address the rotation variation, we jointly learn the optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution. Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation methods, achieving superior accuracy and strong cross-architecture generalization.
@@ 1 Introduction
With the increasing demand for large-scale training datasets, the cost of model retraining from scratch has become a major challenge leading to the research of dataset distillation [21]. The objective of dataset distillation is to produce a significantly smaller synthetic dataset from a large original dataset, that preserves the essential task-relevant information contained in the original dataset. Hence the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [3, 6, 8, 11, 29, 28, 30, 17, 26, 5, 9] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [29] minimizes the difference between the gradients computed on the synthetic and original datasets. However, it potentially overlooks long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [3, 6, 8, 11] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset. Note that these methods train the networks while optimizing the synthetic dataset, and significantly increase the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [28, 30, 17, 26, 5] have been introduced. They randomly initialize the networks without training to extract the features, and compare the feature distributions between the original and synthetic datasets.
@@
While dataset distillation has been extensively studied for structured data such as images [29, 3, 28] and texts [10, 12, 13], its application to 3D point clouds remains almost unexplored. PCC [27] simply applied an existing distillation method for images [29] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity. Unlike the structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (indices of points), that makes direct feature comparison incorrect worsening the performance of distribution matching based dataset distillation. We refer to this issue as semantic misalignment. Moreover, the datasets of 3D point clouds also suffers from rotational variation. 3D point clouds are often captured or synthesized under arbitrary poses due to the absence of canonical orientations. Such rotational difference causes the objects of the same class yield different features from one another, significantly increasing the intra-class variability. As a result, it becomes difficult to construct a representative synthetic dataset that faithfully follows the feature distribution of the original dataset.
@@
  • To the best of our knowledge, we are the first to propose a distribution matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
  • We devised the SADM loss on the sorted features in each channel to preserve the semantic alignment between compared 3D objects.
@@ 2 Related Works
3D point clouds are generally unordered exhibiting irregular characteristics, that makes it difficult to apply the convolution operations in deep neural networks commonly used for images. Early approaches [25, 32] convert the point clouds into structured representation of multi-view images or voxel grids to enable the use of conventional deep learning architectures, such as convolutional neural networks (CNNs). VoxelNet [32] voxelizes the point clouds, but it causes increased memory consumption and quantization errors. To directly process point clouds without conversion to intermediate structured representations, PointNet [14] was introduced as a pioneering work that learns the features directly from unordered raw point clouds. However, it suffers from the limitation in capturing local geometric relationships, and hierarchical CNNs have been developed tailored to 3D point clouds. PointNet++[15] captures local geometric structures through iterative sampling and grouping strategies. On the other hand, PointConv [23] generalizes the convolution operations to irregular point clouds by adaptively weighting the neighboring points based on their spatial distribution. Meanwhile, DGCNN [22] constructs dynamic graphs from point clouds to adaptively capture the semantic information. Attempts have been also made to apply the transformer [20] to 3D point clouds processing, where the Point Transformer [31] utilizes the attention mechanism to capture the long-range dependencies.
@@ 2.2 Coreset Selection
Coreset selection is a technique designed to select representative samples from a given dataset, while maintaining the model’s performance even with the sampled data. The random selection method [16] randomly chooses a subset of data samples from the whole dataset. It is simple but suffers from the robustness due to the lack of informative criteria for sampling. The K-center method [18] selects data points iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution. The Herding method [2, 1] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space.
@@ 2.3 Dataset Distillation
Dataset distillation [21] methods can be largely categorized into the gradient matching [29], trajectory matching [3, 6, 8, 11], and distribution matching [28, 30, 17, 26, 5] approaches. The gradient matching, first introduced by DC [29], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset. The trajectory matching methods [3, 6, 8, 11], initially proposed by MTT [3], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients. ATT [11] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching. The distribution matching, introduced by DM [28], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets. DataDAM [17] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead. Furthermore, M3D [26] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions. Recently, a gradient matching-based point clouds distillation method [27] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
@@ 3.1 Preliminaries
Problem Definition. The objective of dataset distillation is to compress the task-relavant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on ${\boldsymbol{\mathcal{S}}}$ achieves close performance to that trained on ${\boldsymbol{\mathcal{T}}}$. Given a 3D point cloud sample ${{\bf p}}$ following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset ${\boldsymbol{\mathcal{S}}}^{\star}$ can be obtained via
${\boldsymbol{\mathcal{S}}}^{\star}=\mathop{\arg\min}\limits_{{\boldsymbol{\mathcal{S}}}}\mathbb{E}_{{{\bf p}}}\left[||{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{T}}}}}({{\bf p}}),l)-{\mathcal{L}}(\phi_{{{\boldsymbol{\mathcal{S}}}}}({{\bf p}}),l)||^{2}\right],$ (1)
@@
Distribution Matching. Distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
@@
where $\phi$ is the feature extractor and $D$ denotes a distance function. We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [28]. Also, the Maximum Mean Discrepancy (MMD) [7] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
@@ 3.2 Semantically Aligned Distribution Matching
$\phi({{\bf p}})=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
where $N$ denotes the number of points in ${{\bf p}}$, $C$ is the number of feature channels, and ${\bf f}_{i}=\left\{f_{i,1},f_{i,2},...,f_{i,N}\right\}$ represents the feature vector of the $i$-th channel where $f_{i,j}$ is the feature value of the $j$-th point. We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their sizes, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$. Then we have the set of the sorted feature vectors as
@@
$\tilde{K}({{\boldsymbol{\mathcal{T}}},\boldsymbol{\mathcal{S}}})=\frac{1}{|\boldsymbol{\mathcal{T}}|\cdot|\boldsymbol{\mathcal{S}}|}\sum_{{{{\bf t}}\in\boldsymbol{\mathcal{T}}}}\sum_{{{\bf s}}\in\boldsymbol{\mathcal{S}}}\exp\left(-\frac{||\tilde{\phi}({{\bf t}})-\tilde{\phi}({{\bf s}})||^{2}}{2\sigma}\right).$ (7)
@@
where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the most largest feature in each channel. The SADM loss is then formulated as a weighted sum of the two losses, given by
@@
where $\lambda_{1},\lambda_{2}$ are the weighting parameters. By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering sematic structures of 3D point clouds, thereby improving the performance dataset distillation.
@@ 3.3 Estimation of Optimal Rotations
3D objects usually exhibit different orientations from each other. Therefore, while generating optimal 3D objects in the synthetic dataset in terms of their geometric shapes, we also estimate their optimal rotations best representing various orientations of 3D objects in the original dataset. In practice, we introduce three rotation angles, $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, corresponding to rotations around the $x$-, $y$-, and $z$-axes, respectively. These angles are treated as learnable parameters, allowing the orientation of each synthetic 3D object in ${\boldsymbol{\mathcal{S}}}$ to be adaptively adjusted during dataset distillation. Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$, where $\hat{\boldsymbol{\mathcal{S}}}$ is the synthetic dataset whose 3D objects are rotated toward the estimated optimal orientations.
Note that we optimize the synthetic 3D objects as well as their rotation parameters simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then form $\hat{\boldsymbol{\mathcal{S}}}$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero. For each class, a mini-batch is sampled from $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per class. The corresponding synthetic mini-batch is sampled from $\hat{\boldsymbol{\mathcal{S}}}$, with the batch size determined by the number of point cloud objects per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\hat{\boldsymbol{\mathcal{S}}}})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while aligning their orientations more effectively. The benefit of this joint optimization is justified in the following proposition.
$\min_{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (11)
@@
Proof Sketch. The proof is provided in Appendix A. Allowing optimization over both $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set compared to optimizing only $\boldsymbol{\mathcal{S}}$. The feasible set for the joint optimization includes the feasible set of the original optimization as a subset, since optimizing only $\boldsymbol{\mathcal{S}}$ is equivalent to the special case of the joint optimization with the fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$. Consequently, the minimum loss achieved by jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ must be less than or equal to the minimum loss achieved by optimizing only $\boldsymbol{\mathcal{S}}$. $\square$
Table 1: Comparison of quantitative performance. DC, DM, and the proposed method were initialized with random selection [16] for fair comparison. The best and the second best scores are highlighted in bold and underlined, respectively.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
@@
Random 35.5±$\pm 4.7$ 75.2±$\pm 1.7$ 85.3±$\pm 1.1$ 34.6±$\pm 1.8$ 60.0±$\pm 1.32$ 74.1±$\pm 0.4$ 34.0±$\pm 3.0$ 54.8±$\pm 1.5$ 63.1±$\pm 1.0$ 13.9±$\pm 1.4$ 20.4±$\pm 1.3$ 34.8±$\pm 1.1$
Herding 40.1 ±$\pm 5.2$ 78.0 ±$\pm 1.3$ 86.9±$\pm 0.6$ 54.4 ±$\pm 2.0$ 68.5 ±$\pm 1.0$ 78.8 ±$\pm 0.4$ 49.5 ±$\pm 2.3$ 59.8 ±$\pm 0.9$ 66.9 ±$\pm 0.5$ 15.7 ±$\pm 1.7$ 27.7 ±$\pm 1.4$ 38.7 ±$\pm 1.7$
K-center 40.1 ±$\pm 5.2$ 77.6±$\pm 1.9$ 83.2±$\pm 1.4$ 54.4±$\pm 2.0$ 63.0 ±$\pm 2.7$ 65.3 ±$\pm 1.1$ 49.5 ±$\pm 2.3$ 51.4 ±$\pm 1.7$ 47.8 ±$\pm 0.7$ 15.7 ±$\pm 1.7$ 19.8 ±$\pm 0.8$ 24.0 ±$\pm 1.0$
DM 31.9±$\pm 4.3$ 77.6±$\pm 1.6$ 86.1±$\pm 0.8$ 32.3±$\pm 4.8$ 62.0±$\pm 1.9$ 75.2±$\pm 0.5$ 27.8±$\pm 4.0$ 56.1±$\pm 1.2$ 64.3±$\pm 0.7$ 14.9±$\pm 2.0$ 23.3±$\pm 1.5$ 35.6±$\pm 1.4$
DC 43.9±$\pm 5.5$ 75.0±$\pm 2.4$ 86.1±$\pm 1.1$ 52.0±$\pm 1.7$ 66.6±$\pm 1.3$ 75.6±$\pm 0.5$ 49.8±$\pm 1.3$ 60.1±$\pm 1.4$ 64.6±$\pm 0.7$ 18.6±$\pm\textbf{1.6}$ 24.3±$\pm 2.5$ 35.3±$\pm 1.2$
Ours 44.7 ±$\pm\textbf{6.1}$ 84.4 ±$\pm\textbf{1.2}$ 87.8 ±$\pm\textbf{1.0}$ 55.8 ±$\pm\textbf{1.5}$ 72.1 ±$\pm\textbf{0.8}$ 80.1 ±$\pm\textbf{0.4}$ 50.2 ±$\pm\textbf{1.4}$ 63.7 ±$\pm\textbf{0.7}$ 68.4 ±$\pm\textbf{0.5}$ 17.4 ±$\pm 1.5$ 31.6 ±$\pm\textbf{1.0}$ 43.9 ±$\pm\textbf{1.9}$
Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [15], DGCNN [22], PointConv [23], and PT [31], respectively.
Datasets ModelNet10 [24] ModelNet40 [24] ShapeNet [4] ScanObjectNN [19]
@@
PointNet++ 35.0±$\pm 8.3$ 51.5±$\pm 7.8$ 82.9±$\pm\textbf{1.4}$ 14.1±$\pm 2.7$ 46.9±$\pm 6.0$ 68.5±$\pm\textbf{0.5}$ 20.3±$\pm 3.7$ $\pm 3.4$ 53.2±$\pm\textbf{1.5}$ 15.5±$\pm 3.6$ 11.6±$\pm 2.2$ 22.0±$\pm\textbf{3.9}$
DGCNN 56.6±$\pm 5.3$ 62.6±$\pm 2.2$ 79.0±$\pm\textbf{2.2}$ 34.8±$\pm 3.5$ 52.3±$\pm 2.7$ 66.9±$\pm\textbf{1.1}$ 13.8±$\pm 2.2$ 38.0±$\pm 2.7$ 52.8±$\pm\textbf{1.6}$ 17.7±$\pm 2.4$ 12.5±$\pm 1.3$ 19.0±$\pm\textbf{1.5}$
PointConv 33.4±$\pm 6.2$ 40.0±$\pm 10.5$ 56.6±$\pm\textbf{4.9}$ 20.0±$\pm 5.7$ 37.9±$\pm 3.9$ 51.3±$\pm\textbf{6.2}$ 17.4±$\pm 3.0$ 34.0±$\pm 4.8$ 47.4±$\pm\textbf{2.3}$ 15.2±$\pm 1.7$ 14.2±$\pm 1.8$ 16.8±$\pm\textbf{1.5}$
PT 48.5±$\pm 6.6$ $\pm 1.6$ 77.6±$\pm\textbf{1.6}$ 26.7±$\pm 5.0$ 47.5±$\pm 2.7$ 61.6±$\pm\textbf{1.1}$ 34.2±$\pm 5.3$ 42.6±$\pm 2.5$ 55.5±$\pm\textbf{0.5}$ 17.1±$\pm 0.8$ 14.9±$\pm 0.5$ 21.9±$\pm\textbf{2.7}$

### v3

@@ Abstract
Large-scale datasets are usually required to train deep neural networks; however, they increase computational complexity, hindering practical applications. Recently, dataset distillation for images and texts has attracted considerable attention, as it reduces the original dataset to a small synthetic one to alleviate the computational burden of training while preserving essential task-relevant information. However, dataset distillation for 3D point clouds remains largely unexplored, as point clouds exhibit fundamentally different characteristics from those of images, making this task more challenging. In this paper, we propose a distribution-matching-based distillation framework for 3D point clouds that jointly optimizes the geometric structures and orientations of synthetic 3D objects. To address the semantic misalignment caused by the unordered nature of point clouds, we introduce a Semantically Aligned Distribution Matching (SADM) loss, which is computed on the sorted features within each channel. Moreover, to handle rotational variations, we jointly learn optimal rotation angles while updating the synthetic dataset to better align with the original feature distribution. Extensive experiments on widely used benchmark datasets demonstrate that the proposed method consistently outperforms existing dataset distillation approaches, achieving higher accuracy and strong cross-architecture generalization.
@@ 1 Introduction
With the increasing demand for large-scale training datasets, the high cost of retraining models from scratch has become a major challenge, motivating research on dataset distillation [dd]. The objective of dataset distillation is to produce a significantly smaller synthetic dataset from a large original dataset, that preserves the essential task-relevant information contained in the original dataset. Hence the models trained on the reduced synthetic dataset are encouraged to achieve comparable performance to those trained on the original dataset. Existing dataset distillation methods [mtt1, mtt2, mtt3, att, gm, dm, idm, datadam, m3d, exploit, divers] can be broadly classified into gradient matching, trajectory matching, and distribution matching approaches. The gradient matching and trajectory matching methods aim to ensure that the synthetic dataset produces similar optimization dynamics to the original dataset. To this end, the gradient matching method [gm] minimizes the difference between the gradients computed on the synthetic and original datasets. However, it potentially overlooks long-term dependencies in the training process. Rather than comparing individual gradients, the trajectory matching methods [mtt1, mtt2, mtt3, att] encourage the models trained on the synthetic dataset follow similar optimization trajectories to those trained on the original dataset. Note that these methods train the networks while optimizing the synthetic dataset, and significantly increase the computational complexity. To alleviate the computational burden in dataset distillation, the distribution matching techniques [dm, idm, datadam, m3d, exploit] have been introduced. The networks are randomly initialized and used without training to extract features, after which the feature distributions of the original and synthetic datasets are compared.
@@
While dataset distillation has been extensively studied for structured data such as images [gm, mtt1, dm] and texts [textdd1, textdd2, textdd3], its application to 3D point clouds remains almost unexplored. PCC [pcc] simply applied an existing distillation method for images [gm] to the 3D point clouds without considering inherent characteristics of 3D point clouds, however it still suffers from high computational complexity. Unlike the structured data, 3D point clouds consist of unordered and irregularly distributed points in 3D space. Therefore, semantically similar regions across different 3D models are often associated with inconsistent orders (i.e., point indices), which makes direct feature comparison incorrect, thereby worsening the performance of distribution-matching-based dataset distillation. We refer to this issue as semantic misalignment. Moreover, the datasets of 3D point clouds also suffers from rotational variation. 3D point clouds are often captured or synthesized under arbitrary poses due to the absence of canonical orientations. Such rotational differences cause objects of the same class to produce different features, significantly increasing intra-class variability. As a result, it becomes difficult to construct a representative synthetic dataset that faithfully follows the feature distribution of the original dataset.
@@
  • To the best of our knowledge, we are the first to propose a distribution-matching-based dataset distillation method of 3D point clouds, that jointly optimizes the shapes and orientations of synthetic dataset.
  • We devised the SADM loss, computed on the sorted features within each channel, to preserve semantic alignment between the compared 3D objects.
@@ 2 Related Works
3D point clouds are generally unordered exhibiting irregular characteristics, that makes it difficult to apply the convolution operations in deep neural networks commonly used for images. To process such data, early studies [(multiview)multi, (voxel)voxelnet] converted point clouds into structured representations, such as multi-view images or voxel grids, enabling the use of standard convolutional neural networks (CNN) architectures. However, these conversions introduce quantization errors and increase memory overhead. PointNet [pointnet] addressed this limitation by directly learning the features from unordered points, which serves as the backbone in our dataset distillation framework. Subsequent methods, such as PointNet++ [pointnet2], PointConv [pointconv], and DGCNN [dgcnn], extended PointNet by capturing local geometric relationships through hierarchical architecture. Additionally, attempts have been also made to apply the transformer [transformer] to 3D point clouds processing, where Point Transformer [pointtransformer] utilizes the attention mechanism to capture the long-range dependencies. We adopt these architectures as evaluation networks to assess the cross-architecture generalization capability of our distilled datasets.
@@ 2.2 Coreset Selection
Coreset selection is a technique designed to select representative samples from a given dataset, while maintaining the model’s performance even with the sampled data. The random selection method [herding(icarl)] randomly chooses a subset of data samples from the whole dataset. It is simple but suffers from the robustness due to the lack of informative criteria for sampling. The K-center method [kcenter] selects data samples iteratively that maximize the minimum distance to the set of already selected ones, taking into account the data distribution. The Herding method [herding(e2e), herding(scail)] iteratively takes data points that minimize the discrepancy between the mean embeddings of the selected subset and the entire dataset in the feature space. The coreset selection operates within the space of existing samples, and hence lacks the flexibility to synthesize informative patterns. This limitation naturally leads to the emergence of dataset distillation, which generates synthetic datasets capturing task-specific patterns beyond the original samples.
@@ 2.3 Dataset Distillation
Dataset distillation [dd] methods can be largely categorized into the gradient matching [gm], trajectory matching [mtt1, mtt2, mtt3, att], and distribution matching [dm, idm, datadam, m3d, exploit] approaches. The gradient matching, first introduced by DC [gm], minimizes the difference between the gradients computed from the original and synthetic datasets, respectively, guiding the synthetic dataset to follow the training direction of the original dataset. The trajectory matching methods [mtt1, mtt2, mtt3, att], initially proposed by MTT [mtt1], encourage the models trained on the synthetic dataset to follow the optimization trajectories similar to those trained on the original dataset rather than comparing individual gradients. ATT [att] automatically adjusts the length of the training trajectories between the synthetic and original datasets, enabling more effective and precise matching. The distribution matching, introduced by DM [dm], focuses on minimizing the distance between the feature distributions of the original and synthetic datasets. DataDAM [datadam] enhances the distribution matching by aligning the feature maps using attention mechanism, achieving unbiased feature representation with low computational overhead. Furthermore, M3D [m3d] employs the Gaussian kernel function in the distribution matching loss, enabling the alignment across higher-order statistical characteristics of the feature distributions. Recently, a gradient matching-based point clouds distillation method [pcc] has been introduced, however it simply applied the existing image-based method without considering the unique characteristics of 3D point clouds.
@@ 3.1 Preliminaries
Problem Definition. The objective of dataset distillation is to compress the task-relevant information in the original dataset ${\boldsymbol{\mathcal{T}}}=\left\{{{\bf t}}_{i}\right\}_{i=1}^{|\mathcal{T}|}$ and generate a much smaller synthetic dataset ${\boldsymbol{\mathcal{S}}}=\left\{{{\bf s}}_{i}\right\}_{i=1}^{|\mathcal{S}|}$, where $|{\boldsymbol{\mathcal{S}}}|\ll|{\boldsymbol{\mathcal{T}}}|$, such that a model trained on $\boldsymbol{\mathcal{S}}$} achieves close performance to that trained on $\boldsymbol{\mathcal{T}}$}. Given a 3D point cloud sample $\bf p$}} following a real data distribution with the corresponding class label $l$, the optimal synthetic dataset $\boldsymbol{\mathcal{S}}$$^{\star}$ can be obtained via
[…]
@@
Distribution Matching. The distribution matching (DM) strategy focuses on aligning the feature distributions derived from the original and synthetic datasets, respectively, via
@@
where $\phi$ is the feature extractor and $D$ denotes a distance function. We employ a randomly initialized, untrained network, which has been demonstrated to sufficiently capture the structural information for distribution alignment [dm]. Also, the Maximum Mean Discrepancy (MMD) [mmd] loss $\mathcal{L}_{\text{MMD}}$ is often used as $D$, given by
@@ 3.2 Semantically Aligned Distribution Matching
$\phi({{\bf p}})=\begin{bmatrix}f_{1,1}&f_{2,1}&\cdots&f_{C,1}\\ f_{1,2}&f_{2,2}&\cdots&f_{C,2}\\ \vdots&\vdots&\ddots&\vdots\\ f_{1,N}&f_{2,N}&\cdots&f_{C,N}\end{bmatrix}=\left[{\bf f}_{1},{\bf f}_{2},\dots,{\bf f}_{C}\right],\quad{\bf f}_{i}\in\mathbb{R}^{N},$ (5)
where $N$ denotes the number of points in $\bf p$}}, $C$ is the number of feature channels, and ${\bf f}_{i}=\left\{f_{i,1},f_{i,2},...,f_{i,N}\right\}$ represents the feature vector of the $i$-th channel where $f_{i,j}$ is the feature value of the $j$-th point. We then perform channel-wise sorting to the extracted features of $f_{i,j}$’s in the descending order of their values, and obtain the sorted feature vector $\tilde{{\bf f}}_{i}=\left\{\tilde{f}_{i,1},\tilde{f}_{i,2},...,\tilde{f}_{i,N}\right\}$ such that $\tilde{f}_{i,j}\geq\tilde{f}_{i,j+1}$. Then we have the set of the sorted feature vectors as
@@
[…]
@@
where $\tilde{K}_{\text{top}}(\cdot,\cdot)$ denotes the Gaussian kernel computed with only the largest feature in each channel. The SADM loss is then formulated as a weighted sum of the two losses, given by
@@
where $\lambda_{1},\lambda_{2}$ are the weighting parameters. By using $\mathcal{L}_{\text{SADM}}$ instead of $\mathcal{L}_{\text{MMD}}$, we facilitate reliable matching of feature distributions considering semantic structures of 3D point clouds, thereby improving the performance dataset distillation.
@@ 3.3 Estimation of Optimal Rotations
3D objects usually exhibit different orientations from each other. Therefore, while generating optimal 3D objects in the synthetic dataset in terms of their geometric shapes, we also estimate their optimal rotations best representing various orientations of 3D objects in the original dataset. In practice, we introduce three rotation angles, $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, corresponding to rotations around the $x$-, $y$-, and $z$-axes, respectively. These angles are treated as learnable parameters, allowing the orientation of each synthetic 3D object in $\boldsymbol{\mathcal{S}}$} to be adaptively adjusted during dataset distillation. Therefore, instead of minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}})$ in (10), we minimize $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})})$, where $\mathcal{R}_{\boldsymbol{\theta}}$ denotes the rotation operator according to the rotation parameters $\boldsymbol{\theta}=(\theta_{x},\theta_{y},\theta_{z})$. The overall objective is formally defined as
$\left\{{\boldsymbol{\mathcal{S}}}^{\star},{\boldsymbol{\theta}}^{\star}\right\}=\mathop{\arg\min}_{\left\{{\boldsymbol{\mathcal{S}}},{\boldsymbol{\theta}}\right\}}\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})).$ (11)
Note that we optimize the synthetic 3D objects as well as their rotation parameters simultaneously, during the dataset distillation process. Specifically, at each iteration, we randomly initialize the network parameters and construct the synthetic dataset $\boldsymbol{\mathcal{S}}$ by randomly selecting samples from the original dataset. We then form $\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})$ by rotating the objects in $\boldsymbol{\mathcal{S}}$ according to the angles $\theta_{x}$, $\theta_{y}$, and $\theta_{z}$, initially set to zero. For each class, a mini-batch is sampled from $\boldsymbol{\mathcal{T}}$, with a batch size of 8 per class. The corresponding synthetic mini-batch is sampled from $\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})$, with the batch size determined by the number of point cloud objects per class (PPC). The joint optimization process iteratively updates both the geometric structure of the synthetic dataset and their rotation parameters by minimizing $\mathcal{L}_{\text{SADM}}({\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}})})$. This ensures that the synthetic dataset preserves the geometric characteristics of the original dataset while aligning their orientations more effectively. The benefit of this joint optimization is justified in the following proposition.
Proposition 1. Jointly optimizing the synthetic dataset $\boldsymbol{\mathcal{S}}$ and the rotation parameters $\boldsymbol{\theta}$ guarantees a lower or equal loss to that of optimizing $\boldsymbol{\mathcal{S}}$ alone.
$\min_{\left\{\boldsymbol{\mathcal{S}},\boldsymbol{\theta}\right\}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\mathcal{R}_{\boldsymbol{\theta}}(\boldsymbol{\mathcal{S}}))\leq\min_{\boldsymbol{\mathcal{S}}}\mathcal{L}_{\text{SADM}}(\boldsymbol{\mathcal{T}},\boldsymbol{\mathcal{S}}),$ (12)
@@
Proof Sketch. The proof is provided in Appendix A. Jointly optimizing over $\boldsymbol{\mathcal{S}}$ and $\boldsymbol{\theta}$ enlarges the feasible set, as the original optimization over $\boldsymbol{\mathcal{S}}$ alone is a special case of the joint optimization with fixed $(\theta_{x},\theta_{y},\theta_{z})=(0,0,0)$. $\square$
Table 1: Comparison of quantitative performance. DC, DM, and the proposed method were initialized with random selection [herding(icarl)] for fair comparison. ‘Whole’ refers to the classification accuracy obtained by training the network on the entire original dataset without any distillation. ‘Ratio’ represents the percentage of the size of the distilled dataset compared to that of the original dataset. The best and the second best scores are highlighted in bold and underlined, respectively.
Datasets ModelNet10 [modelnet40] ModelNet40 [modelnet40] ShapeNet [shapenet] ScanObjectNN [scanobjectnn]
@@
[…]
Ours 44.7 $\pm\textbf{6.1}$ 84.4 $\pm\textbf{1.2}$ 87.8 $\pm\textbf{1.0}$ 55.8 $\pm\textbf{1.5}$ 72.1 $\pm\textbf{0.8}$ 80.1 $\pm\textbf{0.4}$ 50.2 $\pm\textbf{1.4}$ 63.7 $\pm\textbf{0.7}$ 68.4 $\pm\textbf{0.5}$ 17.4 $\pm 1.5$ 31.6 $\pm\textbf{1.0}$ 43.9 $\pm\textbf{1.9}$
Table 2: Comparison of cross-architecture generalization performance evaluated on PointNet++ [pointnet2], DGCNN [dgcnn], PointConv [pointconv], and PT [pointtransformer], respectively.
Datasets ModelNet10 [modelnet40] ModelNet40 [modelnet40] ShapeNet [shapenet] ScanObjectNN [scanobjectnn]
@@
[…]